# The origin of anomalous non-linear microwave absorption in Josephson junction qubits: mysterious nature of two level systems or their dynamic interaction?

### v1

@@ IV Conclusions
Second, we attempted to interpret the experiment assuming a hypothetical power law distribution of TLS dipole moments $p$ in the form $F(p)\propto p^{\eta-3}$. This distribution, indeed, results in the desirable loss tangent behavior; yet, the proposed hypothesis needs justification. For a justification, we find on the one hand that the proposed distribution of dipole moments is consistent with the experimental observations in amorphous aluminum oxide films [39]. On the other hand, it is consistent with the expected distribution of coupling strengths (dipole moments in our problem) of low-energy collective excitations in systems with long-range interactions [15; 40; 16]. Thus the anomalous loss tangent field dependence signals that the TLSs could be, indeed, represented by such universal excitations. The relevance of theory requires dominating TLS dipole-dipole interactions over elastic interactions. This requires large absolute values of TLS dipole moments, which is a sign of the electronic nature of TLSs in materials, where the anomalous loss tangent behavior is observed.
These theoretical suggestions need further experimental verifications and theoretical developments. Although we ruled out the interaction with two level fluctuators as a source of anomalous non-linear absorption, the interaction effect can be different if, for instance, the fluctuators are multilevel. However, if the anomalous absorption is due to any dynamic interaction (like the interaction with fluctuators) it should be dramatically sensitive to the temperature. Particularly, it should disappear with decreasing the temperature, since at zero temperatures all dynamic excitations must be frozen out. Consequently, the investigation of loss tangent temperature dependence at low temperatures can test the significance of dynamic TLS interactions. The lack of temperature dependence will rule out any dynamic interaction. We did not find any evidences of temperature dependence of a non-linear absorption in the recent work [24].
@@ Appendix A Definition of the fluctuator induced decoherence rate
where index $i$ enumerates all neighboring fluctuators. The decoherence rate $r$ depends on the spatial positions and tunneling amplitudes of all surrounding fluctuators. We estimate the typical value of this rate evaluating its distribution function $P(r)$. The Fourier transform of this distribution function reads [53; 16] ($n$ is the spatial density of fluctuators, see Eq. (6))

### v2

@@ IV Conclusions
Second, we attempted to interpret the experiment assuming a hypothetical power law distribution of TLS dipole moments $p$ in the form $F(p)\propto p^{\eta-3}$. This distribution, indeed, results in the desirable loss tangent behavior; yet, the proposed hypothesis needs justification. For a justification, we find on the one hand that the proposed distribution of dipole moments is consistent with the experimental observations in amorphous aluminum oxide films [39]. On the other hand, it is consistent with the expected distribution of coupling strengths (dipole moments in our problem) of low-energy collective excitations in systems with long-range interactions [15; 40; 16]. Thus the anomalous loss tangent field dependence signals that the TLSs could be, indeed, represented by such universal excitations. The relevance of theory requires dominating TLS dipole-dipole interactions over elastic interactions. This requires large absolute values of TLS dipole moments, which is a sign of the electronic nature of TLSs in materials, where the anomalous loss tangent behavior is observed. Both comparison of the hypothetical dipole moment distribution with the experiment and the theoretical analysis considering the low energy collective excitations suggest that the power $\eta$ in the dipole moment probability density function is close to zero. This assumption is consistent with the interpolation formula used in Ref. [24] to interpret experimental data.
The proposed theory needs further experimental verifications and theoretical developments. Although we ruled out the interaction with two level fluctuators as a source of anomalous non-linear absorption, the interaction effect can be different if, for instance, the fluctuators are multilevel. However, if the anomalous absorption is due to any dynamic interaction (like the interaction with fluctuators) it should be dramatically sensitive to the temperature. Particularly, it should disappear with decreasing the temperature, since at zero temperatures all dynamic excitations must be frozen out. Consequently, the investigation of loss tangent temperature dependence at low temperatures can test the significance of dynamic TLS interactions. The lack of temperature dependence will rule out any dynamic interaction. We did not find any evidences of temperature dependence of a non-linear absorption in the recent work [24].
@@ Appendix A Definition of the fluctuator induced decoherence rate
where index $i$ enumerates all neighboring fluctuators. The decoherence rate $r$ depends on the spatial positions and tunneling amplitudes of all surrounding fluctuators. We estimate the typical value of this rate evaluating its distribution function $P(r)$. The Fourier transform of this distribution function reads [54; 16] ($n$ is the spatial density of fluctuators, see Eq. (6))